# Dynamic Approximate Maximum Matching in the Distributed Vertex Partition Model

### v2

@@ 4.1. Description and Analysis of the Algorithm
    • $w_{i}$ is matched to some node $x\in X^{\prime}$, whereas $u_{i}$ is matched to some node in $V$. Moreover, if $x\neq x_{i}$, i.e., $w_{i}$ was not matched to the same node that it was matched to in $M_{X,W}$, then $x_{i}$ is not free in $M$.
@@
For our analysis, we define graph $G_{*}=G[X\cup W\cup U\cup V]$. Note that $G_{*}\supseteq G_{*}^{\prime}$, and $G_{*}$ may in fact contain up to $\Theta\left(n\right)$ nodes.
@@
The matching $M$ obtained at the end of Phase 3 is maximal and does not admit any 3-augmenting paths in $G_{*}$. In addition, all vertices that were matched at the end of Phase 1 remain matched.
@@
Assume towards a contradiction that there is such a 3-augmenting path $p$. We distinguish several cases, depending on the matched edge $e$ that $p$ traverses.
@@
Consider Case (1), i.e., $p=(a,w_{i},u_{i},b)$ such that $\{w_{i},u_{i}\}\in I$. Since $a$ and $b$ are free, Lemma 12 tells us that $a,b\in G_{*}$. Moreover, by Lemmas 11(B) and 7, we know that $a\in X$ and $b\in V$. Since Stage 2 removes all 3-augmenting paths across $\{w_{i},u_{i}\}$ in the graph $G_{*}^{\prime}$, but $p$ exists, it follows that $a\in X\setminus X^{\prime}$. Moreover, we know that there does not exist any edge $\{x_{i},w_{i}\}$ in the virtual matching $M_{X,W}$, as we did not match $w_{i}$ to $x_{i}$ in Stage 1, which means that $w_{i}$ was unmatched in $M_{X,W}$. Thus, it is possible to add the edge $\{a,w_{i}\}$ to $M_{X,W}$, which, however, contradicts the maximality of $M_{X,W}$ guaranteed by Lemma 9.
Next, we consider Case (2), i.e., the matched edge of $p$ is $e=\{x,w_{i}\}$, where $x\in X$ and $w_{i}\in W$. Then, by Lemmas 6 and 7, it follows that the first (free) node in $p$ must be in $W$, which, however contradicts Lemma 11(A). To complete the proof, observe that the argument for Case (3) is analogous. ∎
@@
Assume towards a contradiction that there exists a 3-augmenting path $p=(a,b,c,d)$ in $G$. By combining Lemmas 13 and 12, we know that the matched edge $\{b,c\}$ cannot be in $G_{*}$, and in particular $\{b,c\}\notin I$, which means that it is not possible for $\{b,c\}$ to be incident to exactly one new edge connecting to a free node. Clearly, $\{a,b\}$ and $\{c,d\}$ cannot both be free edges, as otherwise, we would have resolved the resulting 3-augmenting path already in Phase 1. Thus, the only choice is that $\{a,b\}$ and $\{c,d\}$ are both old edges. Since matched nodes are not unmatched at any point in the algorithm (see Lem. 4(a) and 11(A)), it follows that path $p$ already must have existed at the start of Phase 1, contradicting the assumption that the previous matching did not have any 3-augmenting paths. ∎
@@
To complete the proof of Theorem 2, recall that the running time in Lemma 15 holds for a single mini-batch of $\Theta\left(\sqrt{k\beta}\right)$ edge insertions. By taking a union bound over the $\Theta\left(\ell/\sqrt{k\beta}\right)$ mini-batches, it follows that Lemmas 14 and 15 also hold for the entire batch $B$, and we obtain the claimed update time of $O\left(\frac{\ell\log n}{\sqrt{k\beta}}\right)$.
@@ 4.2. The Lower Bound under Edge Insertions
In the proof of Theorem 16, we extend the lower bound graph that was used in the proof of Theorem 1.1 in (Li et al., 2024). Consider a graph $G$ that consists of $n/5$ segments $L_{1},\ldots,L_{q}$, whereby each $L_{i}$ consists of five nodes $t_{i},u_{i},v_{i},w_{i},x_{i}$ and will contain two edges at first: $\{u_{i},v_{i}\}$ and $\{v_{i},w_{i}\}$. [^5] For simplicity, we assume that $n/5$ and $n/k$ are integers, as this does not impact the asymptotic bounds. We say that $v_{i}$ is the middle vertex of $L_{i}$. If we conceptually place the line segments on a $5\times(n/5)$ grid, we can tile the grid using $k$ ($5\times$)-rectangles $R_{1},\ldots,R_{k}$, to cover all nodes (without overlaps), whereby each rectangle covers exactly $n/k$ vertices. That is, each rectangle represents the subset of vertices hosted by a player. In more detail, we restrict the choice of the rectangles such that every player holds at most one vertex from $\{u_{i},v_{i},w_{i}\}$, for every $i\in[n/5]$. Note that the vertex assignment is possible since we are assuming an adaptive adversary.
The adversary first inserts the edges $\{u_{i},v_{i}\}$ and $\{v_{i},w_{i}\}$ sequentially, for every $i\in[n/5]$. It is easy to see that in the resulting graph, also denoted by $G$, any maximal matching is also a maximum matching. Moreover, there is a player $P$ who hosts a set $S_{P}$ of at least $n/(5k)$ middle vertices $v_{i}$. Notice that the adaptive adversary knows $P$ and $S_{P}$ when selecting the edge insertions. Let $I_{P}$ be the set of indexes of the line segments, for which there is a node $v_{i}\in S_{P}$, and note that $\lvert I_{P}\rvert\geqslant n/(5k)$.
The adversary samples a subset $J_{P}\subseteq I_{P}$ of size $\ell$ u.a.r., and also samples u.a.r. an $\ell$-length vector $\Gamma$ of random bits. Then, for each $i\in J_{P}$, inserts either the edge $\{t_{i},u_{i}\}$ if $\Gamma[i]=0$ (Case 1) or, if $\Gamma[i]=1$, it inserts the edge $\{w_{i},x_{i}\}$ (Case 2). The intuition behind the hardness of the lower bound is that player $P$ hosting the set $S_{P}$ has no way of distinguishing between Case 1 and Case 2 without further communication with the other players. Observe that, before inserting the $\ell$ randomly chose edges in the line segments $L_{i}$ $(i\in[J_{P}])$, either $\{u_{i},v_{i}\}$ or $\{v_{i},w_{i}\}$ must be in the matching. Without loss of generality, assume that $\{u_{i},v_{i}\}$ is matched. Since the player $P$ does not host any vertex incident to an inserted edge, it remains unaware of the edge insertions unless it communicates with other players. It is easy to verify that, the status of the edges $\{u_{i},v_{i}\}$ and $\{v_{i},w_{i}\}$ must change (from matched to unmatched and vice versa) to avoid a 3-augmenting path, if and only if the adversary chooses to insert the edge $\{t_{i},u_{i}\}$ for line segment $L_{i}$, which happens with probability $\frac{1}{2}$.
@@
Let random variable $\mathsf{Out}$ be the edges that are output by $P$ as a part of a maximal matching without 3-augmenting paths after the algorithm’s response to the batch of $\ell$ edge insertions. We use $\textbf{1}_{\text{Succ}}$ to denote the indicator random variable that the algorithm succeeds, and recall that it is $1$ with high probability (i.e., at least 11/n1-1/n).
@@
Now, we can show the lower bound. Let $\Pi$ be the transcript of messages received by player $P$ during the edge insertions. We should notice that $\mathsf{Out}$ is a function of $\Pi$ and the local state of $P$ including the public randomness. Given $P$, we have a Markov Chain: $(P,J_{P})\to\Pi\to\mathsf{Out}$. Then by the data-processing inequality, Lemma 17 and (5), it follows that

### v3

@@ 4.1. Description and Analysis of the Algorithm
    • $w_{i}$ is matched to some node $x\in X^{\prime}$, whereas $u_{i}$ is matched to some node in $V$. Moreover, if $x\neq x_{i}$, i.e., $w_{i}$ was not matched to the same node that it was matched to in the virtual matching $M_{X,W}$, then $x_{i}$ is not free in $M$.
@@
The matching $M$ obtained at the end of Phase 3 is maximal in $G_{*}$ and does not admit any 3-augmenting paths in $G_{*}$. In addition, all vertices that were matched at the end of Phase 1 remain matched.
@@
Assume towards a contradiction that there is a 3-augmenting path $p$. We distinguish several cases, depending on the matched edge $e$ that $p$ traverses.
@@
Consider Case (1), i.e., $p=(a,w_{i},u_{i},b)$ such that $\{w_{i},u_{i}\}\in I$. Since $a$ and $b$ are free, Lemma 11 tells us that $a,b\in G_{*}$. Moreover, by Lemmas 10(B) and 6, we know that $a\in X$ and $b\in V$. Since Stage 2 removes all 3-augmenting paths across $\{w_{i},u_{i}\}$ in the graph $G_{*}^{\prime}$, but $p$ exists, it follows that $a\in X\setminus X^{\prime}$. Moreover, we know that there does not exist any edge $\{x_{i},w_{i}\}$ in the virtual matching $M_{X,W}$, as we did not match $w_{i}$ to $x_{i}$ in Stage 1, which means that $w_{i}$ was unmatched in $M_{X,W}$. Thus, it is possible to add the edge $\{a,w_{i}\}$ to $M_{X,W}$, which, however, contradicts the maximality of $M_{X,W}$ guaranteed by Lemma 8.
Next, we consider Case (2), i.e., the matched edge of $p$ is $e=\{x,w_{i}\}$, where $x\in X$ and $w_{i}\in W$. Then, by Lemmas 5 and 6, it follows that the first (free) node in $p$ must be in $W$, which, however contradicts Lemma 10(A), since all nodes in $W$ were matched. To complete the proof, observe that the argument for Case (3) is analogous. ∎
@@
Assume towards a contradiction that there exists a 3-augmenting path $p=(a,b,c,d)$ in $G$. By combining Lemmas 12 and 11, we know that the matched edge $\{b,c\}$ cannot be in $G_{*}$, and in particular $\{b,c\}\notin I$, which means that it is not possible for $\{b,c\}$ to be incident to exactly one new edge connecting to a free node. Clearly, $\{a,b\}$ and $\{c,d\}$ cannot both be free edges, as otherwise, we would have resolved the resulting 3-augmenting path already in Phase 1. Thus, the only choice is that $\{a,b\}$ and $\{c,d\}$ are both old edges. Since matched nodes are not unmatched at any point in the algorithm (see Lem. 3(a) and 10(A)), it follows that path $p$ already must have existed at the start of Phase 1, contradicting the assumption that the previous matching did not have any 3-augmenting paths. ∎
@@
To complete the proof of Theorem 1, recall that the running time in Lemma 14 holds for a single mini-batch of $\Theta\left(\sqrt{k\beta}\right)$ edge insertions. By taking a union bound over the $\Theta\left(\ell/\sqrt{k\beta}\right)$ mini-batches, it follows that Lemmas 13 and 14 also hold for the entire batch $B$, and we obtain the claimed update time of $O\left(\frac{\ell\log n}{\sqrt{k\beta}}\right)$.
@@ 4.2. The Lower Bound under Edge Insertions
We use a similar lower bound construction as in the proof of Theorem 1.1 of (li2024dynamic). Consider a graph $G$ that consists of $n/5$ segments $L_{1},\ldots,L_{q}$, whereby each $L_{i}$ consists of five nodes $t_{i},u_{i},v_{i},w_{i},x_{i}$ and will contain two edges at first: $\{u_{i},v_{i}\}$ and $\{v_{i},w_{i}\}$. [^4] For simplicity, we assume that $n/5$ and $n/k$ are integers, as this does not impact the asymptotic bounds. We say that $v_{i}$ is the middle vertex of $L_{i}$. If we conceptually place the line segments on a $5\times(n/5)$ grid, we can tile the grid using $k$ ($1\times(n/5k)$)-rectangles $R_{1},\ldots,R_{k}$, to cover all nodes (without overlaps), whereby each rectangle covers exactly $n/5k$ vertices. In more detail, we restrict the choice of the rectangles such that every player holds the vertices of exactly $5$ of these rectangles with the restrictions that the player obtains at most one vertex from $\{u_{i},v_{i},w_{i}\}$, for each index $i\in[n/5]$. Note that the vertex assignment is possible since we are assuming an adaptive adversary.
The adversary first inserts the edges $\{u_{i},v_{i}\}$ and $\{v_{i},w_{i}\}$ sequentially, for every $i\in[n/5]$. It is easy to see that in the resulting graph, also denoted by $G$, any maximal matching is a maximum matching. Moreover, there is a player $P$ who hosts a set $S_{P}$ of at least $n/(5k)$ middle vertices $v_{i}$. Notice that the adaptive adversary knows $P$ and $S_{P}$ when selecting the edge insertions. Let $I_{P}$ be the set of indexes of the line segments, for which there is a node $v_{i}\in S_{P}$, and note that $\lvert I_{P}\rvert\geqslant n/(5k)$.
The adversary samples a subset $J_{P}\subseteq I_{P}$ of size $\ell$ u.a.r., and also samples u.a.r. an $\ell$-length vector $\Gamma$ of random bits. Then, for each $i\in J_{P}$, the adversary inserts either the edge $\{t_{i},u_{i}\}$ if $\Gamma[i]=0$ (Case 1) or, if $\Gamma[i]=1$, it inserts the edge $\{w_{i},x_{i}\}$ (Case 2). The intuition behind the hardness of the lower bound is that player $P$ hosting the set $S_{P}$ has no way of distinguishing between Case 1 and Case 2 without further communication with the other players. Observe that, before inserting the $\ell$ randomly chose edges in the line segments $L_{i}$ $(i\in[J_{P}])$, either $\{u_{i},v_{i}\}$ or $\{v_{i},w_{i}\}$ must be in the matching. Without loss of generality, assume that $\{u_{i},v_{i}\}$ is matched. Since the player $P$ does not host any vertex incident to an inserted edge, it remains unaware of the edge insertions unless it communicates with other players. It is easy to verify that, the status of the edges $\{u_{i},v_{i}\}$ and $\{v_{i},w_{i}\}$ must change (from matched to unmatched and vice versa) to avoid a 3-augmenting path, if and only if the adversary chooses to insert the edge $\{t_{i},u_{i}\}$ for line segment $L_{i}$, which happens with probability $\frac{1}{2}$.
@@
Let random variable $\mathsf{Out}$ be the edges that are output by $P$ as a part of a matching without 3-augmenting paths after the algorithm’s response to the batch of $\ell$ edge insertions. We use $\textbf{1}_{\text{Succ}}$ to denote the indicator random variable that the algorithm succeeds, and recall that it is $1$ with high probability (i.e., with probability at least 11/n1-1/n).
@@
Now, we can show the lower bound. Let $\Pi$ be the transcript of messages received by player $P$ during the edge insertions. We should notice that $\mathsf{Out}$ is a function of $\Pi$ and the local state of $P$ including the public randomness. Given $P$, we have a Markov Chain: $(P,J_{P})\to\Pi\to\mathsf{Out}$. By the data-processing inequality, Lemma 16 and (5), it follows that